# Bayesian Optimisation with Unknown Hyperparameters: Regret Bounds Logarithmically Closer to Optimal

### v1

@@ Appendix A General Hyperparameter case
$\sum_{t\notin\mathcal{T}}r_{t}=\mathcal{O}\left(|A|B+\left(R^{u^{*}}(|S_{t}^{u^{*}}|)+\sqrt{T\xi_{T}}\right)\left(\sqrt{\sum_{u\in U}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}+|A|\right)\right)$
Appendix B Proof of Lemma 3.1
@@
$\frac{R^{\hat{\theta}}(T)}{R^{\theta^{\star}}(T)}=\frac{C\sqrt{T\gamma_{T}^{\hat{\theta}}}\left(\sqrt{\gamma_{T}^{\hat{\theta}}}+B(\hat{\theta},N^{\star})\right)}{C\sqrt{T\gamma_{T}^{\theta^{\star}}}\left(\sqrt{\gamma_{T}^{\theta^{\star}}}+B(\theta^{\star},N^{\star})\right)}\leq 2\left(\frac{\theta^{\star}}{\hat{\theta}}\right)^{\hat{d}}\leq 2\left(\frac{q(i^{\star})}{q(i^{\star}+1)}\right)^{\hat{d}}=2\left(e^{-1/\hat{d}}\right)^{\hat{d}}=\mathcal{O}(1),$
where $i^{\star}=\max\{i\in\mathbb{N}\mid q(i)\geq\theta^{\star}\}$ and as such we have $\theta^{\star}\leq q(i^{\star})$ and $q(i^{\star}+1)\leq\hat{\theta}$. ∎
Appendix C Proof of Theorem 4.1
@@
We start with a similar regret decomposition as in the proof of Theorem 1 in [6]. Let $t_{0}$ be the first iteration, where $\theta^{\star}$ enters the hyperparameter set $\Theta_{t_{0}}$. Before this happens, all hyperparameters in the set are misspecified and as such we cannot guarantee anything about the regret of those iterations. As such, we bound their regret by $2B$, which is the highest possible regret one can suffer at one iteration. We thus get:
$R_{T}=\sum_{t=1,\dots,t_{0}}r_{t}+\sum_{t=t_{0}+1,\dots,T}r_{t}\leq t_{0}2B+\tilde{R}_{T}$
We note that due to how we add new length scales, we have that $t_{0}=g^{-1}(\frac{\theta_{0}}{\theta^{\star}e^{1/\hat{d}}})$ and we defined $\tilde{R}_{T}$ be the cumulative regret of all iterations after $t_{0}$. Let us define the set $\mathcal{T}$ to be the set of all iterations, where at least one hyperparameter was eliminated. We thus get the following regret bound:
$\tilde{R}_{T}=\sum_{t\in\mathcal{T}}r_{t}+\sum_{t\notin\mathcal{T}}r_{t}\leq|\mathcal{T}|B+\sum_{t\notin\mathcal{T}}r_{t}\leq q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)B+\sum_{t\notin\mathcal{T}}r_{t},$
where the last inequality comes from the fact that we cannot reject more hyperparameters than we have considered in total. We now rely on Lemma A.1. In the notation of the lemma, we can write $A=\underset{t=1,\dots,T}{\bigcup}\Theta_{t}$ to mean the set of all length scale values introduced over the course of the algorithm running. We observe that $|A|\leq q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)$ and our optimal base learner is $u^{\star}=\hat{\theta}$. This gives us:
$\sum_{t\notin\mathcal{T}}r_{t}\leq\sum_{t\notin\mathcal{T}}r_{t}=\mathcal{O}\left(q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)B+\left(R^{\hat{\theta}}(T)+\sqrt{T\xi_{T}}\right)\left(\sqrt{\sum_{u\in A}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}+q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)\right)\right)$
To finish the proof we rely on the following Lemma:
Lemma C.1.
For the same type of kernels with different length scales, we have that $\sqrt{\frac{|S_{t}^{\theta}|}{|S_{t}^{\theta^{*}}|}}\leq\left(\frac{\theta_{0}}{\theta^{*}}\right)^{\hat{d}}$, where $\hat{d}=d$ in case of an RBF kernel and $\hat{d}=d+2\nu$ in case of an $\nu$-Matern kernel.
@@
If $|S_{t}^{\theta^{*}}|\geq|S_{t}^{\theta}|$, the bound holds trivially. Thus we will assume $|S_{t}^{\theta^{*}}|<|S_{t}^{\theta}|$, which implies $\gamma_{|S_{t}^{\theta^{*}}|}\geq\gamma_{|S_{t}^{\theta}|}$ and $\theta\geq\theta^{\star}$. The suspected regret bounds are of the form:
$R^{\theta}(t)=C\sqrt{T\gamma_{T}^{\theta}}\left(\sqrt{\gamma_{T}^{\theta}}+B^{\theta}\right).$
Due to the regret balancing condition (Lemma 5.2 of [29]), we must have:
$R^{\theta}(|S_{t}^{\theta}|)\leq 2R^{\theta^{*}}(|S_{t}^{\theta^{*}}|)$
$\sqrt{\frac{|S_{t}^{\theta}|}{|S_{t}^{\theta^{*}}|}}\leq 2\frac{\sqrt{\gamma_{|S_{t}^{\theta^{*}}|}^{\theta}}\left(\sqrt{\gamma_{|S_{t}^{\theta^{*}}|}^{\theta}}+B(\theta^{\star},N^{\star})\right)}{\sqrt{\gamma_{|S_{t}^{\theta}|}^{\theta}}\left(\sqrt{\gamma_{|S_{t}^{\theta}|}^{\theta}}+B^{\theta}\right)}$
Case 1: Consider the case when $\sqrt{\gamma_{|S_{t}^{\theta^{*}}|}^{\theta}}\geq B(\theta^{\star},N^{\star})$. We then have:
$\sqrt{\frac{|S_{t}^{\theta}|}{|S_{t}^{\theta^{*}}|}}\leq 2\frac{\gamma_{|S_{t}^{\theta^{*}}|}^{\theta}}{\gamma_{|S_{t}^{\theta}|}^{\theta}}\leq 2\left(\frac{\theta}{\theta^{\star}}\right)^{\hat{d}}\frac{\gamma_{|S_{t}^{\theta}|}^{\theta}}{\gamma_{|S_{t}^{\theta}|}^{\theta}}=2\left(\frac{\theta}{\theta^{\star}}\right)^{\hat{d}}.$
Case 2 Consider the case when $\sqrt{\gamma_{|S_{t}^{\theta^{*}}|}^{\theta}}<B(\theta^{\star},N^{\star})$. We then have:
$\sqrt{\frac{|S_{t}^{\theta}|}{|S_{t}^{\theta^{*}}|}}\leq 2\frac{\sqrt{\gamma_{|S_{t}^{\theta^{*}}|}^{\theta}}B(\theta^{\star},N^{\star})}{\sqrt{\gamma_{|S_{t}^{\theta}|}^{\theta}}B(\theta,N^{\star})}\leq 2\left(\frac{\theta}{\theta^{\star}}\right)^{\hat{d}/2}\left(\frac{\theta}{\theta^{\star}}\right)^{d/2}\sqrt{\frac{\gamma_{|S_{t}^{\theta}|}^{\theta}}{\gamma_{|S_{t}^{\theta}|}^{\theta}}}\frac{B(\theta,N^{\star})}{B(\theta,N^{\star})}\leq 2\left(\frac{\theta}{\theta^{\star}}\right)^{\hat{d}}.$
We thus get the following final bound on $R_{T}$:
$\mathcal{O}\left(\left(g^{-1}\left(\frac{\theta_{0}}{\theta^{\star}}\right)+\hat{d}\ln g(T)\right)B+\left(R^{\hat{\theta}}(T)+\sqrt{T\xi_{T}}\right)\left(\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{\hat{d}}\sqrt{\hat{d}\ln g(T)}+\hat{d}\ln g(T)\right)\right).$
By Lemma 3.1 we know we can just replace $R^{\hat{\theta}}$ with $R^{\theta^{\star}}$ in the bound above, which finishes the proof.
Corollary C.1.
Under the same conditions as in Theorem 4.1, with probability at least $1-\delta$, Algorithm 1 achieves the following simple regret bound:
$\min_{t=1,\dots,T}r_{t}=\mathcal{O}\left(\left(g^{-1}\left(\frac{\theta_{0}}{\theta^{\star}}\right)+\hat{d}\ln g(T)\right)\frac{B}{T}+\left(r^{\theta^{*}}(T)+\sqrt{\frac{\xi_{T}}{T}}\right)\left(\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{\hat{d}}+\hat{d}\ln g(T)\right)\right),$
where $\hat{d}=d$ in case of an RBF kernel and $\hat{d}=d+2\nu$ in case of a $\nu$-Matérn kernel.
Appendix D Unknown RKHS norm
Within this section, we show how our algorithm can be extended to handle the case of an unknown RKHS norm. Let us define the following candidate-suggesting function for RKHS norm hyperparameter.
Definition D.1.

### v3

@@ Appendix B General Hyperparameter case
$\frac{1}{|S_{t^{\prime}}^{u}|}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{t^{\prime}}^{u}\end{subarray}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)-\frac{2\sum_{t\in S_{t^{\prime}}^{u}}\beta_{t}^{u}\sigma_{t}^{u}(\bm{x}_{t})}{|S_{t^{\prime}}^{u}|}-2\sqrt{\frac{\xi_{t}}{|S_{t^{\prime}}^{u}|}}$
$\leq\frac{1}{|S_{t^{\prime}}^{u^{\star}}|}\sum_{t\in S_{t^{\prime}}^{u^{\star}}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)+2\sqrt{\frac{\xi_{t}}{|S_{t^{\prime}}^{u^{\star}}|}}.$
Rearranging and observing that $r_{t}=f(\bm{x}^{\star})-f(\bm{x}_{t})$, we obtain:
$\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{t^{\prime}}^{u}\end{subarray}}r_{t}\leq$ $2\sum_{t\in S_{t^{\prime}}^{u}}\beta_{t}^{u}\sigma_{t}^{u}(\bm{x}_{t})+2\sqrt{|S_{t^{\prime}}^{u}|\xi_{t}}+\frac{|S_{t^{\prime}}^{u}|}{|S_{t^{\prime}}^{u^{\star}}|}\sum_{t\in S_{t^{\prime}}^{u^{\star}}}r_{t}+2|S_{t^{\prime}}^{u}|\sqrt{\frac{\xi_{t}}{|S_{t^{\prime}}^{u^{\star}}|}}$
$\leq$ $CR^{u}(|S_{t^{\prime}}^{u}|)+2\sqrt{|S_{t^{\prime}}^{u}|\xi_{t}}+\frac{|S_{t^{\prime}}^{u}|}{|S_{t^{\prime}}^{u^{\star}}|}CR^{u^{\star}}(|S_{t^{\prime}}^{u^{\star}}|)+2|S_{t^{\prime}}^{u}|\sqrt{\frac{\xi_{t}}{|S_{t^{\prime}}^{u^{\star}}|}},$
where we use the fact that $2\sum_{t\in S_{t^{\prime}}^{u}}\beta_{t}^{u}\sigma_{t}^{u}(\bm{x}_{t})\leq CR^{u}(S_{t^{\prime}}^{u})$ for some constant $C>0$, which follows from the proof of Theorem 2 of [11] on which we rely to obtain the suspected regret bounds. We also used the fact that due to $u^{\star}$ being well-specified we have that $\sum_{t\in S_{T}^{u^{\star}}}r_{t}\leq CR^{u^{\star}}(|S_{T}^{u^{\star}}|)$. We now apply Lemma B.3 to get that $R^{u}(|S_{t}^{u}|)\leq R^{u^{\star}}(|S_{t}^{u^{\star}}|)+2B^{u^{\star}}$. Substituting that to the bound developed above and noting that by definition $S_{t^{\prime}}^{u}=S_{T}^{u}$, we get:
$\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{T}^{u}\end{subarray}}r_{t}=\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{t^{\prime}}^{u}\end{subarray}}r_{t}$ $\leq\left(\frac{|S_{t^{\prime}}^{u}|}{|S_{t^{\prime}}^{u^{*}}|}+1\right)CR^{u^{\star}}(|S_{t^{\prime}}^{u^{*}}|)+2CB^{u^{\star}}+2\left(\sqrt{\frac{|S_{t^{\prime}}^{u}|}{|S_{t^{\prime}}^{u^{*}}|}}+1\right)\sqrt{|S_{t^{\prime}}^{u}|\xi_{t}}$
$=\left(\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}+1\right)CR^{u^{\star}}(|S_{T}^{u^{*}}|)+2CB^{u^{\star}}+2\left(\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}+1\right)\sqrt{|S_{T}^{u}|\xi_{t}},$
which finishes the proof. ∎
We are now ready to prove the final result of this Section.
Lemma B.5.
Let us run Algorithm 2 for $T$ iterations with a given choice of the hyperparameter-proposing function $a:\mathbb{N}\to\mathcal{U}$. Let $\mathcal{T}$ be the set of all iterations after $t_{0}$, where at least one hyperparameter is rejected by operation in line 11. If we set $\xi_{t}=2\sigma_{N}^{2}\log\frac{|A|\pi^{2}t^{2}}{6\delta_{B}}$ and $\beta^{u}_{t}=B^{u}+\sigma_{N}\sqrt{2(\gamma^{u}_{t-1}+1+\ln(1/\delta_{A}))}$, we then have that with probability as least $1-\delta_{A}-\delta_{B}$:
$\sum_{t\notin\mathcal{T}}r_{t}=\mathcal{O}\left(|A|B^{u^{\star}}+\left(R^{u^{*}}(T)+\sqrt{T\xi_{T}}\right)\left(\sum_{u\in\mathcal{M}_{0}}\sqrt{\frac{|S_{t}^{u}|}{|S_{t^{\prime}}^{u^{\star}}|}}+|A|\right)\right).$
Proof.
We will prove the bound assuming the events of Theorem 2.1 and Lemma B.1 hold. As this happens with probability at least $1-\delta_{A}-\delta_{B}$, the bound also holds with the same probability. We have that:
$\sum_{t\notin\mathcal{T}}r_{t}$ $\leq\sum_{u\in\mathcal{W}}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{T}^{u}\end{subarray}}r_{t}+\sum_{u\in\mathcal{M}_{0}}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{T}^{u}\end{subarray}}r_{t}$
$\leq\sum_{u\in\mathcal{W}}CR^{u}(|S_{T}^{u}|)+\sum_{u\in\mathcal{M}_{0}}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{T}^{u}\end{subarray}}r_{t}$
$\leq\sum_{u\in\mathcal{W}}\left(CR^{u^{\star}}(|S_{T}^{u}|)+2CB^{u^{\star}}\right)+\sum_{u\in\mathcal{M}_{0}}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{T}^{u}\end{subarray}}r_{t}$
$\leq|\mathcal{W}|\left(CR^{u^{\star}}(T)+2CB^{u^{\star}}\right)+\sum_{u\in\mathcal{M}_{0}}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{T}^{u}\end{subarray}}r_{t},$ (3)
where the first transition is due to all hyperparameters in $\mathcal{W}$ being well-specified and the second transition is due to Lemma B.3. We now tackle the second term:
$\sum_{u\in\mathcal{M}_{0}}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{T}^{u}\end{subarray}}r_{t}\leq\sum_{u\in\mathcal{M}_{0}}\left(\left(\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}+1\right)CR^{u^{\star}}(|S_{T}^{u^{*}}|)+2CB^{u^{\star}}+2\left(\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}+1\right)\sqrt{|S_{T}^{u}|\xi_{t}}\right)$
$\leq\left(\sum_{u\in\mathcal{M}_{0}}\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}+|\mathcal{M}_{0}|\right)CR^{u^{\star}}(|S_{T}^{u^{*}}|)+2C|\mathcal{M}_{0}|B^{u^{\star}}+2\left(\sum_{u\in\mathcal{M}_{0}}\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}+|\mathcal{M}_{0}|\right)\sqrt{T\xi_{t}}$
$=\mathcal{O}\left(\left(\sum_{u\in\mathcal{M}_{0}}\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}+|\mathcal{M}_{0}|\right)R^{u^{\star}}(|S_{T}^{u^{*}}|)+|\mathcal{M}_{0}|B^{u^{\star}}+\left(\sum_{u\in\mathcal{M}_{0}}\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}+|\mathcal{M}_{0}|\right)\sqrt{T\xi_{t}}\right),$
where we used Lemma B.4, Cauchy-Schwarz inequality and the fact that $|S_{T}^{u}|\leq T$ for all $u\in A$. Observe, that suspected regret bounds in BO will be of form $R^{u}(T)=\sqrt{T\beta^{u}_{T}\gamma_{T}^{u}}(\sqrt{\gamma_{T}^{u}}+\sqrt{B^{u}})$. Substituting this fact, we get:
$R^{u^{*}}(|S_{T}^{u^{*}}|)\sum_{u\in\mathcal{M}_{0}}\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}$ $\leq\mathcal{O}\left(\sum_{u\in\mathcal{M}_{0}}\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}\sqrt{|S_{T}^{u^{*}}|\beta^{u^{\star}}_{T}\gamma_{T}^{u}}\left(\sqrt{\gamma_{T}^{u^{\star}}}+\sqrt{B^{u^{\star}}}\right)\right)$
$=\mathcal{O}\left(\sum_{u\in\mathcal{M}_{0}}\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}\sqrt{|S_{T}^{u}|\beta^{u^{\star}}_{T}\gamma_{T}^{u}}\left(\sqrt{\gamma_{T}^{u^{\star}}}+\sqrt{B^{u^{\star}}}\right)\right)$
$=\mathcal{O}\left(\sum_{u\in\mathcal{M}_{0}}\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}R^{u^{\star}}(|S_{T}^{u}|)\right)$
$\leq\mathcal{O}\left(\sum_{u\in\mathcal{M}_{0}}\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}R^{u^{\star}}(T)\right).$
We thus get :
$\sum_{u\in\mathcal{M}_{0}}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{T}^{u}\end{subarray}}r_{t}=\mathcal{O}\left(|\mathcal{M}_{0}|B^{u^{\star}}+\left(R^{u^{*}}(T)+\sqrt{T\xi_{T}}\right)\left(\sum_{u\in\mathcal{M}_{0}}\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}+|\mathcal{M}_{0}|\right)\right).$
Substituting this back into Equation 3 yields the following bound:
$\sum_{t\notin\mathcal{T}}r_{t}\leq\mathcal{O}\left(|A|B^{u^{\star}}+\left(R^{u^{*}}(T)+\sqrt{T\xi_{T}}\right)\left(\sum_{u\in\mathcal{M}_{0}}\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}+|A|\right)\right)$
Appendix C Proof of Lemma 3.1
@@
$\frac{R^{\hat{\theta}}(T)}{R^{\theta^{\star}}(T)}=\frac{\sqrt{T\gamma_{T}^{\hat{\theta}}}\left(\sqrt{\gamma_{T}^{\hat{\theta}}}+B(\hat{\theta},N^{\star})\right)}{\sqrt{T\gamma_{T}^{\theta^{\star}}}\left(\sqrt{\gamma_{T}^{\theta^{\star}}}+B(\theta^{\star},N^{\star})\right)}=2\left(\frac{\theta^{\star}}{\hat{\theta}}\right)^{d}\leq 2\left(\frac{q(i^{\star})}{q(i^{\star}+1)}\right)^{d}=2\left(e^{-1/d}\right)^{d}=\mathcal{O}(1),$
where $i^{\star}=\max\{i\in\mathbb{N}\mid q(i)\geq\theta^{\star}\}$ and as such we have $q(i^{\star}+1)=\hat{\theta}\leq\theta^{\star}\leq q(i^{\star})$. ∎
Appendix D Proof of Theorem 4.1
@@
We start with a similar regret decomposition as in the proof of Theorem 1 in [6]. Let $t_{0}$ be the first iteration, where a length scale value smaller or equal to $\theta^{\star}$ enters the hyperparameter set $\Theta_{t_{0}}$. We will refer to that value as $\hat{\theta}$. Before this happens, all hyperparameters in the set are misspecified and as such we cannot guarantee anything about the regret of those iterations. As such, we bound their regret by $2B^{\star}$, which is the highest possible regret one can suffer at one iteration. We thus get:
$R_{T}=\sum_{t=1,\dots,t_{0}}r_{t}+\sum_{t=t_{0}+1,\dots,T}r_{t}\leq t_{0}2B^{\star}+\tilde{R}_{T}$ (4)
We note that due to how we add new length scales, we have that $t_{0}\leq g^{-1}(\frac{\theta_{0}}{\theta^{\star}e^{1/d}})$ and we defined $\tilde{R}_{T}$ be the cumulative regret of all iterations after $t_{0}$. Let us define the set $\mathcal{T}$ to be the set of all iterations, where at least one hyperparameter was eliminated. We thus get the following regret bound:
$\tilde{R}_{T}=\sum_{t\in\mathcal{T}}r_{t}+\sum_{t\notin\mathcal{T}}r_{t}\leq 2|\mathcal{T}|B^{\star}+\sum_{t\notin\mathcal{T}}r_{t}\leq 2q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)B^{\star}+\sum_{t\notin\mathcal{T}}r_{t},$
where the last inequality comes from the fact that we cannot reject more hyperparameters than we have considered in total. We now rely on Lemma B.5, which provides a bound on $\sum_{t\notin\mathcal{T}}r_{t}$ with probability at least $1-\delta_{A}-\delta_{B}$ and we set $\delta_{A}=\delta_{B}=\delta/2$. In the notation of the Lemma, we can write $A=\underset{t=1,\dots,T}{\bigcup}\Theta_{t}$ to mean the set of all length scale values introduced over the course of the algorithm running and by $\mathcal{M}_{0}$ we mean all length scales longer than $\hat{\theta}$ that were selected at least once after $t_{0}$. We observe that $|A|\leq q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)$ and our optimal base learner is $u^{\star}=\hat{\theta}$. This gives us:
$\sum_{t\notin\mathcal{T}}r_{t}=\mathcal{O}\left(q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)B^{\star}+\left(R^{\hat{\theta}}(T)+\sqrt{T\xi_{T}}\right)\left(\sum_{u\in\mathcal{M}_{0}}\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}+q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)\right)\right)$
To finish the proof we rely on the following Lemma, which we prove in Appendix E.
Lemma D.1.
If the event of Lemma B.5 holds, then for any $\theta\in\mathcal{M}_{0}$ and $t\geq t_{0}$ we have that $\sqrt{\frac{|S_{t}^{\theta}|}{|S_{t}^{\theta^{\star}}|}}\leq\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}$.
We thus get the following final bound:
$\sum_{t\notin\mathcal{T}}r_{t}=\mathcal{O}\left(q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)B^{\star}+\left(R^{\hat{\theta}}(T)+\sqrt{T\xi_{T}}\right)\left(|\mathcal{M}_{0}|\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}+q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)\right)\right).$
We now observe that $|\mathcal{M}_{0}|=\mathcal{O}\left(q^{-1}\left(\theta^{\star}\right)\right)=\mathcal{O}\left(d\ln\frac{\theta_{0}}{\theta^{\star}}\right)$. We substitute the bound above to Equation 4, together with bound on $|\mathcal{M}_{0}|$ to obtain:
$R_{T}\leq\mathcal{O}\left(\left(g^{-1}\left(\frac{\theta_{0}}{\theta^{\star}e^{1/d}}\right)+\iota\right)B^{\star}+\left(R^{\hat{\theta}}(T)+\sqrt{T\xi_{T}}\right)\left(\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}d\ln\theta^{\star}+\iota\right)\right),$
where $\iota=d\ln g(T)$. By Lemma 3.1 we know we can just replace $R^{\hat{\theta}}(T)$ with $R^{\theta^{\star}}(T)$ in the bound above, which finishes the proof.
Appendix E Proof of Lemma D.1
See D.1
@@
If $|S_{t}^{\theta^{\star}}|\geq|S_{t}^{\theta}|$, the bound holds trivially. Thus we will assume $|S_{t}^{\theta^{\star}}|<|S_{t}^{\theta}|$. The suspected regret bounds are of the form:
$R^{\theta}(T)=\sqrt{T\gamma_{T}^{\theta}}\left(\sqrt{\gamma_{T}^{\theta}}+B(\theta,N^{\star})\right).$
If the event on Lemma B.5 holds that means the event of Lemma B.3 holds as well. Due to the regret balancing condition from Lemma B.3 and the non-triviality of bounds, we have:
$R^{\theta}(|S_{t}^{\theta}|)\leq R^{\theta^{\star}}(|S_{t}^{\theta^{\star}}|)+2B^{u^{\star}}\leq 2R^{\theta^{\star}}(|S_{t}^{\theta^{\star}}|)$
$\sqrt{\frac{|S_{t}^{\theta}|}{|S_{t}^{\theta^{\star}}|}}\leq 2\frac{\sqrt{\gamma_{|S_{t}^{\theta^{\star}}|}^{\theta}}\left(\sqrt{\gamma_{|S_{t}^{\theta^{\star}}|}^{\theta}}+B(\theta^{\star},N^{\star})\right)}{\sqrt{\gamma_{|S_{t}^{\theta}|}^{\theta}}\left(\sqrt{\gamma_{|S_{t}^{\theta}|}^{\theta}}+B(\theta,N^{\star})\right)}.$
To finish the proof we consider the following two cases.
Case 1: Consider the case when $\sqrt{\gamma_{|S_{t}^{\theta^{\star}}|}^{\theta}}\geq B(\theta^{\star},N^{\star})$. We then have:
$\sqrt{\frac{|S_{t}^{\theta}|}{|S_{t}^{\theta^{\star}}|}}\leq 2\frac{\gamma_{|S_{t}^{\theta^{\star}}|}^{\theta}}{\gamma_{|S_{t}^{\theta}|}^{\theta}}\leq 2\left(\frac{\theta}{\theta^{\star}}\right)^{d}\frac{\gamma_{|S_{t}^{\theta}|}^{\theta}}{\gamma_{|S_{t}^{\theta}|}^{\theta}}=2\left(\frac{\theta}{\theta^{\star}}\right)^{d}\leq 2\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}.$
Case 2 Consider the case when $\sqrt{\gamma_{|S_{t}^{\theta^{\star}}|}^{\theta}}<B(\theta^{\star},N^{\star})$. We then have:
$\sqrt{\frac{|S_{t}^{\theta}|}{|S_{t}^{\theta^{\star}}|}}$ $\leq 2\frac{\sqrt{\gamma_{|S_{t}^{\theta^{\star}}|}^{\theta}}B(\theta^{\star},N^{\star})}{\sqrt{\gamma_{|S_{t}^{\theta}|}^{\theta}}B(\theta,N^{\star})}\leq 2\left(\frac{\theta}{\theta^{\star}}\right)^{d/2}\sqrt{\frac{\gamma_{|S_{t}^{\theta}|}^{\theta}}{\gamma_{|S_{t}^{\theta}|}^{\theta}}}\left(\frac{\theta}{\theta^{\star}}\right)^{d/2}\frac{B(\theta,N^{\star})}{B(\theta,N^{\star})}$
$\leq 2\left(\frac{\theta}{\theta^{\star}}\right)^{d}\leq 2\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}.$
Appendix F Unknown RKHS norm
Within this section, we show how our algorithm can be extended to handle the case of an unknown RKHS norm. Let us define the following candidate-suggesting function for the RKHS norm hyperparameter.
Definition F.1.